The recommended daily value (RDV) of calcium for men and women between the ages 19 – 50 is 1000 mg. Powder milk contains 33.0% of the RDV per serving. The mass of a serving of powder milk is 24 g. What is the (wt/wt %) of calcium per serving of powder milk?
Answer #1

total mass of milk powder 24 g

let the mass of calcium be x gm in the milk powder

percentage of calcium in the milk powder P = x/24 *100

given P =33

33=x/24*100

33/100 =x/24

.33*24 =x

7.92 g= x

7.92 gis the (wt/wt %) of calcium per serving of powder milk?